OpenStudy (anonymous):

Find an equivalent fraction for the decimal number. In your final answer, include all of your work. 0.61

OpenStudy (anonymous):

to convert a decimal to a fraction multiply and divide with 100 in this case because 0.61 has 2 decimal \[0.61 \times \frac{ 100 }{ 100}=\frac{ 61 }{ 100 }\] see if you can still simplify it(it isnt possible in this case ) and thats you fraction